# Seeing What’s Not There: Spurious Correlation in Multimodal LLMs

### v1

@@ Appendix E Object Detector Validation
Without Spur. Cue: “street sign”
@@
Without Spur. Cue: “bird”
@@
Without Spur. Cue: “goggles”
@@
Without Spur. Cue: “glass”
@@
Often, a majority of images in the dataset have 0 $f_{i}$-scores for a given spurious feature $f_{i}$. To avoid bias from the initial dataset ordering, we randomize the order of 0-score images before producing the ranking.
To demonstrate the quality of object detections that we used, we randomly selected some COCO classes, and for these, uniformly sampled some images from the top $K=100$ and bottom $K=100$ of one of the rankings considered by SpurLens. The results are shown in Fig. 10.
Appendix F Token Dropping
F.1 Implementation
@@
Suppose that we have an image of a target object, and have a bitmask of the object for that image. We aim to drop visual tokens from Qwen2-VL’s processing of the image, without affecting its vision of other areas. After some basic resizing, Qwen2-VL breaks the image into patches of size $14\times 14$ pixels. However, adjacent $2\times 2$ tokens are later merged into a single token by a simple MLP layer [32]. Therefore, when dropping tokens, we must effectively work with a “token size” of $28\times 28$ pixels.
@@
Figure 11 illustrates the procedure using an image/mask pair from the “howler monkey” class of HardImageNet. Figure 11 (a) and (b) show the image and mask, and (c) shows the regions of the image that would be dropped, which is a superset of the mask. (Note that we visualize this as a mask, but the model would not see black masked-out regions – the tokens for those regions would not be present at all.)
F.2 Validation
We devise a small demonstration to show the validity of our method: we pass images to Qwen2-VL with the prompt “Is something wrong with this image?”. The responses with various images from Figure 11 are as follows:
  • With original image (a): “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene.”
  • With the masked image (b): “Yes, there appears to be an issue with the image. The silhouette of the animal seems to be floating in the air without any visible connection to the tree branches or leaves. This suggests that the image might be digitally altered or manipulated, as it is not a natural occurrence for an animal to be floating in the air without support.”
  • With the token-dropped image (visualized as a mask in (c)): “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene.”
  • We add uniform random noise to the area of the image that will be dropped to obtain (d). We then pass in the corrupted image with the same tokens dropped as in (c). The model answers: “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene.”
@@ F.2 Validation
F.3 Comparison with Masking
In Sec. 6.1, we computed HR Gaps on HardImagenet classes, using token-dropping to create artificial samples without the target object. Here, we compare those results to the results if we were to use masking (filling in the masked area with 0 RGB pixel values) to create the artificial negative samples.
@@ F.3 Comparison with Masking
Appendix G Spurious Bias in MLLMs: Evidence Without SpurLens
In this section, we provide evidence that spurious bias in MLLMs can be observed even without using SpurLens. Instead of systematically identifying high-spurious images, we leverage existing benchmarks, Spurious ImageNet [20] and COCO [9], to demonstrate this issue.
Spurious ImageNet is a subset of 100 classes from ImageNet [25], containing 75 images per class that include spurious cues but lack the target object. For baseline images, we randomly selected 75 images from ImageNet for each class.
@@ Appendix G Spurious Bias in MLLMs: Evidence Without SpurLens
We illustrate some hallucinations produced by GPT-4o-mini in this setup in Fig. 13. The quantitative results for this experiment are shown in Table 6. For COCO, even with our simple setup for identifying spurious images, the HR is at least three times higher than that of the baseline images. For Spurious ImageNet, which systematically identifies spurious cues for each class, the ratios are even larger. Additionally, we visualized the HR distribution across different classes in Fig. 12. Our findings align with [19], highlighting that the bias from reliance on spurious cues is critically class-dependent.
@@
Appendix H Vision Encoder Experiments
@@
After training the logistic regression classifier on this dataset, we estimate the PA $\text{PA Gap}=\text{PA}_{s}-\text{PA}_{c}$ as before by computing the accuracy on the $K=100$ most spurious and $K=100$ least spurious images in HardImageNet-train of class $c$ (note that these were excluded from the training dataset). Additionally, we run the classifier on all of HardImageNet-val to get a better understanding of its overall accuracy.
@@
Figure 15: PA Gaps separately for each HardImageNet class for all experiments in Fig. 14. The colors for different values of $f$ are the same as in that figure.
@@
Appendix I SpurLens Validation
@@
We compare the PA Gaps of the spurious ranking selected by SpurLens to the PA Gap (with $K=50$) obtained with the original HardImageNet spurious ranking [18]. Additionally, we compare it against a random baseline constructed as follows. For each (class, model), we take 16 random rankings, compute the PA Gap, and take the maximum; this procedure is then performed 16 times, we take the average as the baseline PA Gap for that (class, model) pair. Comparisons against these baselines provide evidence that we are capturing a statistically significant signal.
Next, we apply SpurLens to 79 COCO classes. (We exclude the ”person“ class because (1) it is extremely generic and broad, making it outstandingly difficult to find strong spurious features for, and (2) it is several times larger than the second-largest class, making it very computationally expensive.) We use GPT-4 to generate 32 potential spurious features, and go through the same SpurLens process to compute PA Gaps with $K=100$. Likewise, we compute random baselines for each class in the same manner.
@@
Finally, we apply SpurLens to the 100 Imagenet classes from SpuriousImagenet [20]. PA Gaps are computed with $K=50$, but otherwise the same procedure is used as for COCO.
@@
Qwen2-VL 5.9 5.8 5.3 12.9 8.7 5.9 5.0
Llama-3.2 13.7 12.1 8.5 12.9 8.6 16.7 9.4
LLaVA-v1.6 12.4 7.5 7.4 13.7 8.4 14.9 8.4
GPT-4o-mini 12.6 6.3 9.1 17.9 10.6 11.3 8.3
@@
The class-wise averaged results of these experiments are presented in Table 7. We note that, while strong spurious features are not present in every class, SpurLens does tend to find significant spurious cues for most classes. The spurious gaps found by SpurLens for each class can be found in Appendix L for HardImageNet, Appendix M for COCO, and Appendix N for the ImageNet subset. These results provide evidence that our methodology is both statistically significant and more effective than past work.
Appendix J Class-wise Plots
We provided plots for other models in Fig. 16.
@@
Figure 16: Class-wise distribution of PA and HR for experiments in Sec. 5.1 and Sec. 5.2 respectively. These highlight the class-dependent nature of spurious biases.
Appendix K Qualitative Examples of Spuriousity Gap
We highlighted some failures identified by SpurLens in Fig. 17.
@@
With Cue Refer to caption Prompt: Determine whether there is a bus in the image.
Qwen2VL: Yes, there is a bus in the image.
Without Cue Refer to caption Prompt: Determine whether there is a bus in the image.
Qwen2VL: No, there is no bus in the image.
@@
With Cue Refer to caption Prompt: Do you see a tennis racket in the image?
Qwen2VL: Yes, there is a tennis racket in the image.
Without Cue Refer to caption Prompt: Do you see a tennis racket in the image?
Qwen2VL: No, there is no tennis racket in the image.
@@
With Cue Refer to caption Prompt: Is there a hot dog in the image?
LLaVA-v1.6: Yes, there is a hot dog in the image. It is placed in a bun and topped with macaroni and cheese, and it appears to be covered in a sauce or condiment.
Without Cue Refer to caption Prompt: Is there a hot dog in the image?
LLaVA-v1.6: No, there is no hot dog in the image. The image shows a variety of sandwiches and pastries displayed in a bakery or deli case.
@@
With Cue Refer to caption Prompt: Is there a cake in the image?
Qwen2VL: Yes, there is a cake in the image.
Without Cue Refer to caption Prompt: Is there a cake in the image?
Qwen2VL: There is no existence of a cake in the image description.
@@
With Cue Refer to caption Prompt: Determine whether there is a vase in the image.
Llama-3.2: Yes, there is a vase in the image. The vase is blue and contains orange flowers.
Without Cue Refer to caption Prompt: Determine whether there is a vase in the image.
Llama-3.2: There are no vases in the image. The objects in the image appear to be candlesticks.
@@
Appendix L Class-wise SpurLens Results on HardImageNet
As described in Sec. 4, for each class, GPT-4 is used to generate 32 potential spurious features. We compute the PA Gap with $K=50$ for these features, and choose feature with the maximum Gap. The results are in Table 8.
@@
0 dog sled sky 2.0 mountain 16.0 collar 10.0 first aid kit 19.3
1 howler monkey fallen leaves 0.7 fruit 12.0 sky 4.0 fruit 6.7
@@
3 ski slope 2.7 slope 4.7 tree 4.0 goggles 6.0
4 sunglasses water bottle 2.0 beach chair 8.0 pool chair 2.0 camera 6.0
@@
7 gymnastic horizontal bar scoreboard 10.0 scoreboard 14.0 water bottle 5.3 rope 10.7
8 patio beverage 6.7 umbrella 9.3 rug 5.3 beverage 8.7
@@
10 miniskirt grass 10.7 grass 28.7 street sign 32.0 crop top 10.0
11 keyboard space bar desk 2.7 desk 28.0 cable 11.3 screen 42.0
@@
1 howler monkey fallen logs 8.7 fallen leaves 2.7 fallen leaves 9.3
2 seat belt headrest 12.7 headrest 27.3 headrest 38.0
@@
10 miniskirt sidewalk 41.3 handbag 8.0 street sign 36.7
@@
Appendix M Class-wise SpurLens Results on COCO
As described in Sec. 4, for 79 COCO classes, we use GPT-4 to generate 32 potential spurious features. We exclude COCO class 1 “person” from our analysis because (1) it is extremely generic and broad, making it outstandingly difficult to find good spurious features for, and (2) it is several times larger than the second-largest class, making it computationally expensive to analyze.

### v2

@@ F.1 Agreement with Object Detector
Without Spur. Cue: “street sign"
@@
Without Spur. Cue: “bird"
@@
Without Spur. Cue: “goggles"
@@
Without Spur. Cue: “glass"
@@
To provide qualitative support, we also visualize example detections. We randomly selected 4 COCO classes and uniformly sampled images from the top- and bottom-ranked sets (based on cue presence) from one of the SpurLens rankings. Example detections are shown in Figure 9.
It is worth noting that we also experimented with using Grounding DINO (liu2024groundingdinomarryingdino) as an open-set object detector in place of OWLv2. However, we found that its detections were qualitatively less consistent and reliable. As a result, we use OWLv2 exclusively in our pipeline.
F.2 Agreement on Spurious Gap
In the second study, we assess whether the gaps computed by SpurLens align with those derived from human-annotated spurious cues. We randomly sampled 20 object-cue pairs from COCO and asked human annotators to label whether the associated spurious feature was present in each image. Using these annotations, we recompute the PA Spurious Gap with $K=50$, and compare it to the gap produced by SpurLens using the same $K$. As shown in Figure 10, the two sets of gap values are highly correlated (Pearson $r=0.988$), supporting the claim that SpurLens produces human-aligned and reliable rankings.
Figure 10: Comparison of SpurLens PA Gap and human-labeled PA Gap (both computed with $K=50$) across 20 COCO classes. The strong correlation (Pearson $r=0.988$) indicates that SpurLens produces rankings consistent with human judgments.
Appendix G CelebA Evaluation
While SpurLens focuses on spurious biases related to object recognition and hallucination, prior work has extensively studied other forms of bias, such as correlations between gender and blond hair in the CelebA (celeba) dataset, particularly in unimodal classification settings (dac; groupdro; lastlayer). Following this line of work, we examine whether such spurious correlations persist in MLLMs.
We adopt a similar evaluation setup: using the prompt “Does the person in the image have blond hair? Answer with ‘Yes’ or ‘No’,” we measure accuracy across groups with and without the spurious feature (gender). Results are shown in Table 7. While Spurious Gap still exists, we argue that it is not particularly meaningful in this context. In prior work, Unimodal classifiers trained with ERM performed poorly on the worst group (e.g., blond males), often below 50%. In contrast, MLLMs like Qwen2-VL achieve over 96% accuracy on this group in a zero-shot setting. Moreover, the CelebA annotations are noisy, and in many cases where Qwen’s predictions differ from the labels, the true hair color is ambiguous even to humans. Figure 11 shows qualitative examples.
Table 7: Qwen2-VL accuracy on CelebA across gender and hair color. The spurious feature in this setting is gender. All numbers are percentages.
Hair Color Male Female Gap
Blond 96.63 98.28 1.65
Non-blond 81.70 74.84 6.86
Figure 11: Qualitative examples from CelebA where Qwen-VL’s answer (“Not blond”) differs from the annotated ground truth (“Blond”). Hair color is often ambiguous, even for humans.
Appendix H Waterbirds Evaluation
Similar to CelebA, the Waterbirds dataset has often been used to study the effect of spurious correlations in vision classifiers. Unlike CelebA, however, it is a synthetic dataset with a simplified structure. This may pose issues for MLLMs, which may err due to the synthetic, out-of-distribution nature of the inputs rather than the spurious correlations. As our goal is to draw conclusion about the performance of modern MLLMs in practical, real-world settings, the results of such an analysis may not be fully applicable.
Additionally, Waterbirds was designed for evaluating spurious correlations in a training setup. In the training phase, the dataset creates strong spurious cues (e.g., landbirds predominantly being on land and waterbirds on water). This creates a high spurious correlation that the model learns to rely on. However, during testing, the dataset balances these cues, which causes a drop in classifier accuracy. In contrast, our work focuses on a zero-shot setup, where the model is not explicitly trained on such cues, and any spurious correlations that arise are more naturally occurring. Thus, we believe evaluating on such a synthetic, domain-specific setup would not align with the goal of evaluating MLLMs in more complex, real-world scenarios.
Nevertheless, we still explore Waterbirds’ applicability within the context of MLLMs. We use the following prompt: “We call a bird ’waterbird’ if it is a seabird (albatross, auklet, cormorant, frigatebird, fulmar, gull, jaeger, kittiwake, pelican, puffin, or tern) or waterfowl (gadwall, grebe, mallard, merganser, guillemot, or Pacific loon). Is there a waterbird in the image? Answer with ’Yes’ or ’No’.” This prompt is similar to our standard prompts, prepended by a definition of the target class “waterbird” (following the definition provided by groupdro). The results are presented in Table 8. We observe that all MLLMs examined still exhibited significant spurious gaps, which we interpret as a clear indicator of residual spurious correlations in the models.
Table 8: Accuracy of various open-source models on Waterbirds. All numbers are percentages.
Model Waterbird (Water) Waterbird (Land) Landbird (Water) Landbird (Land)
Qwen 94.58 81.35 64.02 96.03
LLaVA 59.33 13.24 77.38 99.76
Llama 94.34 79.90 38.99 70.88
Appendix I Hyperparameter Sensitivity Analysis
We study how the Spurious Gap varies with the hyperparameter $K$, which controls the number of top- and bottom-ranked images used to compute the gap. As shown in Figure 12, all models exhibit a similar trend: the Spurious Gap decreases for larger values of $K$ and stabilizes as $K$ increases. This behavior is expected: smaller $K$ focuses on more extreme samples and amplifies the gap, while larger $K$ smooths the estimate by averaging over a broader set of images.
Importantly, the gap values begin to stabilize around $K\approx 60$, indicating that SpurLens produces consistent rankings and that our reported gaps are not highly sensitive to this parameter. These trends support the robustness of our method and suggest that our main findings are not artifacts of a specific $K$ choice.
Figure 12: PA (Left) and HR (Right) Spurious Gap as a function of $K$, the number of top/bottom images used in evaluation. Gaps stabilize for $K\sim 60$, suggesting robustness to this parameter.
Appendix J Prompting Mitigation Discussion
This is an extended discussion of the attempted prompt-based mitigation from Section 6.2.
Table 4 includes $\text{HR}_{b}$: the hallucination rates on blank (fully black) images. While $\text{HR}_{b}$ is not always zero in the Baseline setting, it drops to zero with Ensemble or Dual Prompting, suggesting that these strategies help suppress some hallucinations.
The results in Table 4 reveal a trade-off between accuracy and hallucination rates across the reasoning-oriented strategies. Dual Prompting reduces HR, but at the cost of lower PA, while Guiding Prompting improves PA but increases hallucinations, suggesting that encouraging the model to focus on key details might reinforce reliance on spurious cues. In contrast, Prompt Ensembling provides a minor but consistent improvement across all cases, reducing HR while slightly increasing PA, making it the most balanced strategy. None of these increases or decreases are very significant, and none of these strategies resolve the Spurious Gap.
The final two strategies, Spurious List and Spurious Top, aim to leverage the textual information retrieved by the SpurLens pipeline. Since SpurLens leverages LLMs to identify spurious biases in natural language for each class, a natural idea is to provide a list of such potential features to the examined LLMs, which may cause them to recognize and/or avoid the bias; this is the basis of the Spurious List strategy. After running the SpurLens pipeline, we know which of the proposed spurious features is strongest for each class. Providing only this feature (rather than the list of all features, not all of which may be spurious) is not proper, as it contains information from the entire class dataset upon which we are also evaluating. Nevertheless, even with this additional information, we find that spurious bias still clearly persists. In fact, the results for the Spurious List and Spurious Top strategies are often weaker than the reasoning and ensembling-based strategies
The Spurious Gap remains present across all prompting strategies, suggesting that while prompt engineering can influence model behavior, it does not eliminate the underlying reliance on spurious cues. This indicates that the Spurious Gaps detected by SpurLens are a more fundamental issue that cannot be resolved solely through prompting techniques.
Appendix K Token Dropping
K.1 Implementation
@@ Appendix I Hyperparameter Sensitivity Analysis
Suppose that we have an image of a target object, and have a bitmask of the object for that image. We aim to drop visual tokens from Qwen2-VL’s processing of the image, without affecting its vision of other areas. After some basic resizing, Qwen2-VL breaks the image into patches of size $14\times 14$ pixels. However, adjacent $2\times 2$ tokens are later merged into a single token by a simple MLP layer (Qwen2VL). Therefore, when dropping tokens, we must effectively work with a “token size" of $28\times 28$ pixels.
@@
Figure 13 illustrates the procedure using an image/mask pair from the “howler monkey" class of HardImageNet. Figure 13 (a) and (b) show the image and mask, and (c) shows the regions of the image that would be dropped, which is a superset of the mask. (Note that we visualize this as a mask, but the model would not see black masked-out regions – the tokens for those regions would not be present at all.)
K.2 Validation
We devise a small demonstration to show the validity of our method: we pass images to Qwen2-VL with the prompt “Is something wrong with this image?". The responses with various images from Figure 13 are as follows:
  • With original image (a): “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene."
  • With the masked image (b): “Yes, there appears to be an issue with the image. The silhouette of the animal seems to be floating in the air without any visible connection to the tree branches or leaves. This suggests that the image might be digitally altered or manipulated, as it is not a natural occurrence for an animal to be floating in the air without support."
  • With the token-dropped image (visualized as a mask in (c)): “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene."
  • We add uniform random noise to the area of the image that will be dropped to obtain (d). We then pass in the corrupted image with the same tokens dropped as in (c). The model answers: “The image appears to be a close-up of tree branches and leaves. There doesn’t seem to be anything inherently wrong with the image; it’s a natural scene."
@@
K.3 Comparison with Masking
In Section 6.1, we computed HR Gaps on HardImageNet classes, using token-dropping to create artificial samples without the target object. Here, we compare those results to the results if we were to use masking (filling in the masked area with 0 RGB pixel values) to create the artificial negative samples.
@@
Appendix L Spurious Bias in MLLMs: Evidence Without SpurLens
In this section, we provide evidence that spurious bias in MLLMs can be observed even without using SpurLens. Instead of systematically identifying high-spurious images, we leverage existing benchmarks, Spurious ImageNet (spnet) and COCO (coco), to demonstrate this issue.
Spurious ImageNet is a subset of 100 classes from ImageNet (imagenet), containing 75 images per class that include spurious cues but lack the target object. For baseline images, we randomly selected 75 images from ImageNet for each class.
@@
We illustrate some hallucinations produced by GPT-4o-mini in this setup in Figure 15. The quantitative results for this experiment are shown in Table 10. For COCO, even with our simple setup for identifying spurious images, the HR is at least three times higher than that of the baseline images. For Spurious ImageNet, which systematically identifies spurious cues for each class, the ratios are even larger. Additionally, we visualized the HR distribution across different classes in Figure 14. Our findings align with moayeri2023spuriosity, highlighting that the bias from reliance on spurious cues is critically class-dependent.
@@
Appendix M Vision Encoder Experiments
@@
After training the logistic regression classifier on this dataset, we estimate the $\text{PA Gap}=\text{PA}_{s}-\text{PA}_{c}$ as before by computing the accuracy on the $K=100$ most spurious and $K=100$ least spurious images in HardImageNet-train of class $c$ (note that these were excluded from the training dataset). Additionally, we run the classifier on all of HardImageNet-val to get a better understanding of its overall accuracy.
@@
Figure 17: PA Gaps separately for each HardImageNet class for all experiments in Figure 16. The colors for different values of $f$ are the same as in that figure.
@@
Finally, we directly compare the gaps obtained through these vision encoder logistic regressions with the PA gaps from evaluation on Qwen2-VL on the HardImageNet rankings. Using $f=7$, $K=50$, and taking the maximum possible value of $x$ for each dataset (that being $2K$ less than the size of the HardImageNet-train for that class), we obtain a PA Gap on the vision encoder logistic regression classifier for each HardImageNet class. The class-wise average of these gaps is 0.082. Comparing these gaps class-wise against the PA Gaps obtained when evaluating Qwen2-VL on the HardImageNet rankings (with $K=50$), we find a linear correlation coefficient of 0.25. This suggests that the vision encoder is partially responsible for the observed spurious bias in the full MLLM.
Appendix N SpurLens Validation
@@
We compare the PA Gaps of the spurious ranking selected by SpurLens to the PA Gap (with $K=50$) obtained with the original HardImageNet spurious ranking (hardimagenet). Additionally, we compare it against a random baseline constructed as follows. For each (class, model), we take 16 random rankings, compute the PA Gap, and take the maximum; this procedure is then performed 16 times, we take the average as the baseline PA Gap for that (class, model) pair. Comparisons against these baselines provide evidence that we are capturing a statistically significant signal.
Next, we apply SpurLens to 79 COCO classes. (We exclude the "person“ class because (1) it is extremely generic and broad, making it outstandingly difficult to find strong spurious features for, and (2) it is several times larger than the second-largest class, making it very computationally expensive.) We use GPT-4 to generate 32 potential spurious features, and go through the same SpurLens process to compute PA Gaps with $K=100$. Likewise, we compute random baselines for each class in the same manner.
@@
Finally, we apply SpurLens to the 100 Imagenet classes from SpuriousImagenet (spnet). PA Gaps are computed with $K=50$, but otherwise the same procedure is used as for COCO.
@@
Qwen2-VL 5.9 5.8 5.3 15.2 8.7 6.7 5.0
Llama-3.2 15.0 12.1 8.5 14.8 8.6 17.7 9.4
LLaVA-v1.6 13.2 7.5 7.4 16.0 8.4 15.7 8.4
GPT-4o-mini 12.6 6.3 9.1 20.4 10.6 11.9 8.3
@@
The class-wise averaged results of these experiments are presented in Table 11. We note that, while strong spurious features are not present in every class, SpurLens does tend to find significant spurious cues for most classes. The spurious gaps found by SpurLens for each class can be found in Appendix S.1 for HardImageNet, Appendix S.2 for COCO, and Appendix S.3 for the ImageNet subset. These results provide evidence that our methodology is both statistically significant and more effective than past work.
Appendix O Alternative Object Detectors
In the SpurLens pipeline, we use OWLv2 (owlv2) for open-set object detection to identify spurious features in a given image dataset. The reliability of OWLv2 is established in Appendix F. Here, we investigate using other open-set object detectors: GroundingDINO (liu2024groundingdinomarryingdino) and YOLO World (cheng2024yoloworldrealtimeopenvocabularyobject). (Note that closed-vocabulary object detectors, such as DETR and many popular R-CNN variants, are not applicable to our methodology as we need to identify arbitrary spurious features/objects.)
Empirically, we observe that different object detectors have have different confidence behaviors and characteristics. GroundingDINO object detection results tend to have many false positives (incorrect labeled bounding boxes with fairly high confidence). Some qualitative examples of this are provided in Figure 18. This introduces significant noise, which leads to ranking errors and thus errors in Spurious Gap computation, especially for relatively small datasets such as HardImageNet classes ($\sim 1000$ images per class). In contrast, YOLO World tends to be very conservative in its confidence score estimates, and often misses instances of spurious objects; OWLv2 lies in the middle of these extremes, with the best overall performance that we observe empirically.
Figure 18: Qualitative examples comparing object detection with GroundingDino and OWLv2, taken from HardImageNet. OWLv2 is much more conservative with its bounding box proposals, while GroundingDino has many false positives before identifying the most pertinent spurious features.
Figure 19: Plots comparing $f_{i}$-scores from OWLv2 to those from GroundingDINO and YOLO World object detection on two chosen (target object, spurious feature) pairs from HardImageNet. We observe that the correlation between the scores for the object detection models is object-dependent. We also observe the characteristics of these models: GroundingDINO tends to provide high detection scores to many images, while YOLO World is far far more conservative and gives near-0 scores to most images. OWLv2 lies between these extremes.
Nevertheless, we do find some agreement between $f_{i}$-scores computed with these different object detection systems. Figure 19 provides examples of correlations between OWLv2-based $f_{i}$-scores and both GroundingDINO-based $f_{i}$-scores and YOLO World-based $f_{i}$-scores, for select (target object, spurious feature) pairs from HardImageNet. We observe that the correlations are generally strong but varied and class/cue dependent. Nevertheless, they provide us with confidence that our pipeline is sound, and that OWLv2 is a strong choice of object detector for SpurLens.
Appendix P Dataset Diversity Evaluation
In Section 4, SpurLens assumes that we have access to a “large dataset of images”, and that the dataset is sufficiently diverse to support object detection for spurious feature identification. Here, we clarify these heuristics and outline a simple procedure to examine a dataset’s adequacy for use with SpurLens.
A simple minimum requirement for a spurious feature to be “sufficiently represented” in a dataset is for there to be at least $K$ instances where it is present, and at least $K$ instances where it is not, where $K$ is the number of high-spurious and low-spurious images used for the MLLM evaluation. When analyzing GPT4-suggested spurious features, likely some will be sufficiently represented in a dataset, and some will not; a user would likely want at least $\widetilde{N}$ of the features to be sufficiently represented, for some reasonable value of $\widetilde{N}$. We would then run SpurLens on these features, and take the feature with the highest Spurious Gap, which we hope is positive, indicating that we identified a spurious correlation.
Suppose a user has a large dataset, and would like to test its diversity and suitableness before running the MLLM evaluation (the most expensive part of SpurLens). The user could perform feature proposal with GPT-4 to obtain $N$ potential spurious cues (adjusting the prompt and providing additional domain-specific information, as needed, in order to obtain high-quality features for the downstream application). The user would choose $\widetilde{N}\leq N$, the minimum number of sufficiently-represented spurious cues they would like to evaluate on (using their heuristics and domain-specific knowledge to estimate how many spurious cues would likely be needed before getting at least one positive result). The user would perform the object detection step with OWLv2 on all $N$ potential spurious cues.
For a given threshold $\tau\in[0,1]$, a simple criteria is to believe that potential spurious cues with $f_{i}\text{-scores}>\tau$ have the image, and cues with $f_{i}\text{-score}<\tau$ do not have the image. Given $\tau$, a user can compute the maximum value of $K$ for each spurious cue such that it is sufficiently represented in the dataset. Finally, the user can compute the maximum $K$ such that at least $\widetilde{N}$ of the spurious cues are sufficiently represented; call denote value $K_{\tau,\widetilde{N}}$. The user perform this calculation for various values of $\tau$, and choose $\tau^{\star}$ with the maximal $K_{\tau^{\star},\widetilde{N}}$. At this stage, the user can decide if the dataset is sufficiently diverse: if $K_{\tau^{\star},\widetilde{N}}$ is sufficiently large, then they may proceed with the MLLM evaluation; otherwise, they must obtain a new dataset. The chosen value of $K$ determines the accuracy and precision of SpurLens’ PA Gap estimates; the user must determine what value of $K$ is sufficiently large, based on their domain understanding and downstream objective.
As a case study, we apply this method to HardImageNet. In Figure 20, for various values of $\widetilde{N}$, we plot the minimum value of $K_{\tau^{\star},\widetilde{N}}$ over all 15 HardImageNet classes against a range of $\tau$ values. We observe that $K_{\tau^{\star},\widetilde{N}}$ is fairly high (over 100 in all cases). Additionally, if place harse restrictions on object detection confidence (such as $\tau>0.3$), we are still able to achieve large $K$ values ($K>50$). This gives us confidence that HardImageNet is sufficiently diverse, and can support spurious feature detection and evaluation according to the standard SpurLens pipeline.
Figure 20: For various $\widetilde{N}$, we plot the minimum value of $K_{\tau,\widetilde{N}}$ over all 15 HardImageNet class for $\tau\in[0,0.4]$. We observe that peak of each curve is above $K=100$, supporting the fact that HardImageNet classes are diverse datasets and can support spurious object evaluation with (at least) $\widetilde{N}=7$.
Appendix Q Class-wise Plots
We provided plots for other models in Figure 21.
@@
Figure 21: Class-wise distribution of PA and HR for experiments in Section 5.1 and Section 5.2 respectively. These highlight the class-dependent nature of spurious biases.
Appendix R Qualitative Examples of Spurious Gap
We highlighted some failures identified by SpurLens in Figure 22.
@@
With Cue Refer to caption Prompt: Determine whether there is a bus in the image. Qwen2VL: Yes, there is a bus in the image.
Without Cue Refer to caption Prompt: Determine whether there is a bus in the image. Qwen2VL: No, there is no bus in the image.
@@
With Cue Refer to caption Prompt: Do you see a tennis racket in the image? Qwen2VL: Yes, there is a tennis racket in the image.
Without Cue Refer to caption Prompt: Do you see a tennis racket in the image? Qwen2VL: No, there is no tennis racket in the image.
@@
With Cue Refer to caption Prompt: Is there a hot dog in the image? LLaVA-v1.6: Yes, there is a hot dog in the image. It is placed in a bun and topped with macaroni and cheese, and it appears to be covered in a sauce or condiment.
Without Cue Refer to caption Prompt: Is there a hot dog in the image? LLaVA-v1.6: No, there is no hot dog in the image. The image shows a variety of sandwiches and pastries displayed in a bakery or deli case.
@@
With Cue Refer to caption Prompt: Is there a cake in the image? Qwen2VL: Yes, there is a cake in the image.
Without Cue Refer to caption Prompt: Is there a cake in the image? Qwen2VL: There is no existence of a cake in the image description.
@@
With Cue Refer to caption Prompt: Determine whether there is a vase in the image. Llama-3.2: Yes, there is a vase in the image. The vase is blue and contains orange flowers.
Without Cue Refer to caption Prompt: Determine whether there is a vase in the image. Llama-3.2: There are no vases in the image. The objects in the image appear to be candlesticks.
@@
Appendix S Class-wise SpurLens Results
S.1 HardImageNet
As described in Section4, for each class, GPT-4 is used to generate 32 potential spurious features. We compute the PA Gap with $K=50$ for these features, and choose feature with the maximum Gap. The results are in Table 12.
@@
0 dog sled goggles 2.0 snow 22.0 snow 16.7 goggles 14.7
1 howler monkey tree branch 2.0 tree branch 23.3 tree branch 6.7 tree branch 12.0
@@
3 ski helmet 2.0 chairlift 4.0 tree 4.0 goggles 6.0
4 sunglasses water bottle 2.0 water bottle 5.3 camera 2.0 camera 6.0
@@
7 gymnastic horizontal bar scoreboard 10.0 foam pit 17.3 foam pit 8.7 rope 10.7
8 patio table 4.7 umbrella 9.3 rug 5.3 table 8.0
@@
10 miniskirt grass 10.7 grass 28.7 bench 27.3 crop top 10.0
11 keyboard space bar monitor stand 4.0 printer 29.3 monitor stand 14.0 screen 42.0
@@
1 howler monkey fallen logs 8.7 banana 2.7 fallen leaves 9.3
2 seat belt cup holder 12.7 headrest 27.3 headrest 38.0
@@
10 miniskirt sidewalk 41.3 handbag 8.0 sidewalk 36.7
@@
S.2 COCO
As described in Section 4, for 79 COCO classes, we use GPT-4 to generate 32 potential spurious features. We exclude COCO class 1 “person" from our analysis because (1) it is extremely generic and broad, making it outstandingly difficult to find good spurious features for, and (2) it is several times larger than the second-largest class, making it computationally expensive to analyze.